Musich Corporation has an activity-based costing system with three activity cost pools-Machining, Setting Up, and Other. The companys overhead costs, which consist of equipment depreciation and indirect labor, have been allocated to the cost pools already and are provided in the table below. Pool Machining Setting Up Other Total $9200 $ 48,200 $ 23,600$ 81,000 Indirect labor 4200 2800 4000 11,000 Total $ 1340051,000 $ 27.600 92,000 Costs in the Machining cost pool are assigned to products based on machine-hours (MHs) and costs in the Setting Up cost pool are assigned to products based on the number of batches. Costs in the Other cost pool are not assigned to products. Data concerning the two products and the companys costs appear below: MHs Batches Product Z3 Product T1 Total 5900 6100 12.000 700 1300 Product Z3 Product T1 Sales (total) Direct materials (total) Direct labor (total) $ 228,800 s 254,500 S 83,500 S 97,400 S 109,800 S 104,700

Required:

Calculate the following:

Machining Activity Rate

$

Setting up Activity Rate

$

Amount of OH applied to product Z3 (round to the nearest dollar)

$

Amount of OH applied to product T1 (round to the nearest dollar)

$

Product Margin – Z3 (round to the nearest dollar)

$

Product Margin – T1 (round to the nearest dollar)
Answer #1

1) Overhead rate of machining = 13400/12000 = 1.11667 per machine hour

2) Setting up activity rate = 51000/2000 = 25.5 per setup

3) Overhead applied to product Z3 = (13400/12000*5900+51000*700/2000) = $24438

4) Overhead applied to product T1 = (13400/12000*6100+51000*1300/2000) = $39962

5 Product margin Z3 = 228800-83500-109800-24438 = 11062

6) Product margin T1 = 254500-97400-104700-39962 = 12438